An electronics firm is currently manufacturing an item that has a variable cost of $0.50 per unit and a selling price of $1.10 per unit. Fixed costs are $14,000. Current volume is 35,000 units. The firm can substantially improve the product quality by adding a new piece of equipment at an additional fixed cost of $6,400. Variable cost would increase to $0.65, but volume should jump to 50,000 units due to a higher-quality product.

Based on the given information, the decision should be to ____, since the profit with the existing equipment is $__ (enter your response to the nearest whole number and include a minus sign if necessary).
Answer #1

For the existing equipment

  • Variable cost (VC) = 0.50
  • Selling price (SP) = $1.10
  • Fixed cost (FC) = $14000
  • Volume of output (Q) = 35000 units
  • Profit = Q(SP-VC) - FC = 35000(1.10-0.50) - 14000 = (35000 × 0.6) - 14000 = 21000-14000 = $7000

For the new equipment

  • Variable cost (VC) = $0.65
  • Selling price (SP) = $1.10
  • Fixed cost (FC) = $14000+$6400 = $20400
  • Volume of output (Q) = 50000 units
  • Profit = Q(SP-VC) - FC = 50000(1.10-0.65) - 20400 = (50000 × 0.45) - 20400 = 22500-20400 = $2100

Based on the given information the decision should be to choose the existing equipment since the profit with the existing equipment is $7000